# Universal Skeleton Understanding via Differentiable Rendering and MLLMs

### 18 Mar 2026 v1

@@ 1 Introduction
Human skeleton sequences are a compelling case study. Skeletons offer a compact, privacy-preserving, and appearance-invariant representation of motion (Sun et al., 2022). This makes them attractive for applications where raw video is impractical due to bandwidth, storage, or privacy constraints, such as healthcare and human-robot interaction (Wang et al., 2025). However, skeleton data faces a critical format silo problem. Skeleton topologies vary widely across acquisition systems: Kinect v2 provides 25 joints (Zhang, 2012), MoCap systems use 22 SMPL joints (Loper et al., 2015), and 2D pose estimators output 17 COCO keypoints (Sun et al., 2019). This heterogeneity means a model trained on one format cannot be directly applied to another without costly retraining or error-prone joint remapping. Beyond recognition, skeleton-based methods also lack open-ended semantic understanding—they output discrete labels but cannot explain why an action occurs or answer natural language questions about fine-grained body dynamics. These limitations—format silos and shallow semantics—have long hindered the goal of universal skeleton understanding: a single model that comprehends any skeleton format across diverse tasks.
@@ 2.2 LLM/MLLM for Motion Understanding
The rise of Large Language Models (LLMs) and Multimodal LLMs (MLLMs) has opened new avenues for motion understanding. A prominent approach is to tokenize continuous motion into a discrete “pose vocabulary” via VQ-VAE, as in MotionGPT (Jiang et al., 2023) and MotionLLM (Chen et al., 2025). This allows LLMs to process motion as a sequence of tokens. However, the tokenization paradigm introduces critical limitations: (1) the quantization process is inherently lossy, discarding subtle kinematic details; (2) the learned codebook generalizes poorly to novel skeleton formats; and (3) it imposes an artificial semantic gap, forcing the model to learn an abstract “pose language” instead of leveraging its profound, pre-existing visual understanding.
An alternative strategy is to encode skeletons into the embedding space of MLLMs, as explored by SKI models (Sinha et al., 2025). However, this coordinate-based projection struggles to fully activate the MLLM’s powerful pretrained perception of visual patterns and lacks a unified mechanism to natively handle diverse skeleton formats. In contrast, our work introduces a differentiable renderer that translates skeletons into the visual modality that MLLMs natively understand. This allows gradients from the MLLM to directly refine the rendering process, learning a visual representation that is maximally informative for the task at hand.
@@ 2.3 Universal Skeleton Representation
Achieving a universal skeleton representation that transcends format heterogeneity is a long-standing challenge. Prior attempts include joint remapping heuristics, zero-padding to a maximum joint count, or learning format-specific adapters. These solutions are either lossy, introduce noise, or require retraining for each new format. Our approach sidesteps these issues entirely. By rendering any skeleton—regardless of its joint count, connectivity, or coordinate system—into a visually consistent image sequence, DrAction creates a format-agnostic visual lingua franca. This allows a single model to seamlessly process data from Kinect, MoCap, 2D pose estimators, or any other source, enabling true cross-format generalization without architectural modifications or retraining.
@@ 3.2 DrAction: Differentiable Skeleton Renderer
For a skeleton with $J$ joints and varying bone connectivity, we instantiate $K=J+K_{\text{bone}}$ Gaussians. The first $J$ Gaussians are anchored at joints, while $K_{\text{bone}}$ Gaussians are uniformly sampled along bone edges via linear interpolation. This density ensures continuous visual connectivity even for sparse skeletons. Each Gaussian $k$ in canonical space is parameterized as:
@@
Per-Joint Rigid Transforms. For each joint $i$ at frame $t$, we compute a rigid transformation $\mathbf{T}_{i}\in\mathrm{SE}(3)$ that maps from the canonical pose to the current pose. Given the current joint position $\mathbf{j}_{i}^{t}$ and orientation quaternion $\mathbf{q}_{i}^{t}$ (when available), the transformation is:
@@
where $\mathbf{j}_{i}^{c}$ denotes the canonical joint position. When orientation data is unavailable (e.g., for 2D pose estimations), we default to identity rotations, gracefully degrading to translation-only skinning.
Blending and SO(3) Projection. Each Gaussian $k$ is associated with the skeleton via blend weights $\mathbf{w}_{k}\in\Delta^{J-1}$ (on the probability simplex). The blended translation and rotation are computed as:
@@
For each Gaussian $k$, we aggregate position $\mathbf{p}_{k}^{t}=\sum_{i}w_{k,i}\mathbf{j}_{i}^{t}$ and velocity $\mathbf{v}_{k}^{t}=\sum_{i}w_{k,i}\dot{\mathbf{j}}_{i}^{t}$ from associated joints, concatenate with base appearance from an MLP, and optionally process through a GRU for temporal modeling. The NFM then predicts RGB and opacity residuals $(\Delta\text{RGB}_{k},\Delta\alpha_{k})$ plus a saliency gate $g_{k}$. The final color blends the modulated appearance with a depth-based colormap: $\mathbf{C}_{k}=(1{-}\lambda)\sigma(\text{RGB}_{k}^{\text{base}}{+}\Delta\text{RGB}_{k})+\lambda\mathbf{C}_{k}^{\text{depth}}$, where $\lambda$ is learnable.
@@ 3.4 Cooperative Training Strategy
To address “brittle semantics” between similar actions (e.g., “rubbing hands” vs. “clapping”), we introduce a contrastive-like binary classification task. We construct hard negative pairs and ask the model: “Does this video show [Action A]?” (when it actually shows similar Action B). We update the renderer ($\Theta_{\text{render}}$) and projector ($\Theta_{\text{proj}}$), minimizing binary cross-entropy loss $\mathcal{L}_{\text{Disc}}$. This directs the renderer to attend to subtle, discriminative motion details.
@@
Moving beyond recognition to deep understanding, we distill structured reasoning from a stronger teacher model (Hurst et al., 2024). We prompt the teacher to generate a causal chain of thought: step-by-step analysis of body-part movements, temporal evolution, and potential intent, followed by a conclusion. Our model is trained to generate this complete rationale given the rendered video. We optimize the renderer ($\Theta_{\text{render}}$), projector ($\Theta_{\text{proj}}$), and LLM (via LoRA, $\Theta_{\text{LoRA}}$) using auto-regressive loss $\mathcal{L}_{\text{CR}}$. This instills structured, causal action understanding, enabling the model to explain why an action is classified as such.
@@
Finally, we freeze the mature renderer and update only the projector ($\Theta_{\text{proj}}$) and LLM ($\Theta_{\text{LoRA}}$) to refine the mapping for open-vocabulary recognition, minimizing standard cross-entropy loss $\mathcal{L}_{\text{MQA}}$ on target labels.
@@
More design and implementation details on DrAction, Disc-FT, and CR-Distill, as well as ablations on progressive training, are provided in the Appendix.
Table 1: Top-1 accuracy (%) of various methods on NTU-60 and NTU-120. Each split is denoted as X/Y, where X is the number of seen classes and Y is the number of unseen classes. The best results are in red, and the second-best are blue. Results for Qwen2.5-VL-7B and InternVL3-8B were obtained by rendering skeletons with fixed visualization and finetuning on MQA for 6 epochs.
@@ 4.1 Experimental Setup
  • Finetuned MLLMs: Qwen2.5-VL and InternVL3 with fixed skeleton visualization, finetuned on MQA recognition for 6 epochs.
@@ E.2.3 Neural Feature Modulator (NFM) Architecture
An optional single-layer GRU (hidden size 10) models temporal dependencies across frames, producing $\mathbf{h}_{k}^{t}=\text{GRU}(\mathbf{x}_{k}^{t},\mathbf{h}_{k}^{t-1})$. A two-layer MLP ($10\to 64\to 5$) with ReLU activation then predicts:

### 31 Mar 2026 v2

@@ 1 Introduction
Human skeleton sequences are a compelling case study. Skeletons offer a compact, privacy-preserving, and appearance-invariant representation of motion (Sun et al., 2022). This makes them attractive for applications where raw video is impractical due to bandwidth, storage, or privacy constraints, such as healthcare and human-robot interaction (Wang et al., 2025b). However, skeleton data faces a critical format silo problem. Skeleton topologies vary widely across acquisition systems: Kinect v2 provides 25 joints (Zhang, 2012), MoCap systems use 22 SMPL joints (Loper et al., 2015), and 2D pose estimators output 17 COCO keypoints (Sun et al., 2019). This heterogeneity means a model trained on one format cannot be directly applied to another without costly retraining or error-prone joint remapping. Beyond recognition, skeleton-based methods also lack open-ended semantic understanding—they output discrete labels but cannot explain why an action occurs or answer natural language questions about fine-grained body dynamics. These limitations—format silos and shallow semantics—have long hindered the goal of universal skeleton understanding: a single model that comprehends any skeleton format across diverse tasks.
@@ 2.2 LLM/MLLM for Motion Understanding
The rise of Large Language Models (LLMs) and Multimodal LLMs (MLLMs) has opened new avenues for motion understanding. A prominent approach is to tokenize continuous motion into a discrete “pose vocabulary” via VQ-VAE, as in MotionGPT (Jiang et al., 2023) and MotionLLM (Chen et al., 2025). This allows LLMs to process motion as a sequence of tokens. However, the tokenization paradigm introduces critical limitations: (1) the quantization process is inherently lossy, discarding subtle kinematic details; (2) the learned codebook generalizes poorly to novel skeleton formats; and (3) it imposes an artificial semantic gap, forcing the model to learn an abstract “pose language” instead of leveraging its profound, pre-existing visual understanding. SUGAR (Ye et al., 2026) instead adopts a more efficient coordinate-to-language pipeline via visual-motion supervision and Temporal Query Projection. An alternative strategy is to encode skeletons into the embedding space of MLLMs, as explored by SKI models (Sinha et al., 2025). However, this coordinate-based projection struggles to fully activate the MLLM’s powerful pretrained perception of visual patterns and lacks a unified mechanism to natively handle diverse skeleton formats. In contrast, our work introduces a differentiable renderer that translates skeletons into the visual modality that MLLMs natively understand. This allows gradients from the MLLM to directly refine the rendering process, learning a visual representation that is maximally informative for the task at hand.
@@ 2.3 Universal Skeleton Representation
Achieving a universal skeleton representation that transcends format heterogeneity is a long-standing challenge. Prior attempts include joint remapping heuristics (Duan et al., 2022), zero-padding to a maximum joint count (Wang et al., 2024), or learning format-specific adapters (Guo et al., 2023; Wang et al., 2025a). These solutions are either lossy, introduce noise, or require retraining for each new format. Our approach sidesteps these issues entirely. By rendering any skeleton—regardless of its joint count, connectivity, or coordinate system—into a visually consistent image sequence, DrAction creates a format-agnostic visual lingua franca. This allows a single model to seamlessly process data from Kinect, MoCap, 2D pose estimators, or any other source, enabling true cross-format generalization without architectural modifications or retraining.
@@ 3.2 DrAction: Differentiable Skeleton Renderer
For a skeleton with $J$ joints and varying bone connectivity, we instantiate $K=J+K_{\text{bone}}$ Gaussians, where $K_{\text{bone}}=|\mathcal{E}|\times N_{\text{samples}}$ ($|\mathcal{E}|$: number of bone edges from the skeleton’s adjacency list; $N_{\text{samples}}{=}10$: intermediate points per bone). The first $J$ Gaussians are anchored at joints, while $K_{\text{bone}}$ Gaussians are uniformly sampled along bone edges via linear interpolation. Because both $J$ and $|\mathcal{E}|$ are read from the input skeleton, the primitive count adapts automatically to any topology (see Appendix C.2). This density ensures continuous visual connectivity even for sparse skeletons. Each Gaussian $k$ in canonical space is parameterized as:
@@
Per-Joint Rigid Transforms. For each joint $i$ at frame $t$, we compute a rigid transformation $\mathbf{T}_{i}\in\mathrm{SE}(3)$ that maps from the canonical pose to the current pose. While Section 3.1 defines a skeleton purely by joint positions, some capture systems (e.g., Kinect v2) also record per-joint orientation quaternions that describe the local rotation of the bone segment attached to each joint; these rotations are applied to the Gaussian primitives bound to that joint, not to the joint point itself. Given the current joint position $\mathbf{j}_{i}^{t}$ and orientation quaternion $\mathbf{q}_{i}^{t}$ (when available), the transformation is:
@@
where $\mathbf{j}_{i}^{c}$ denotes the canonical joint position—the position of joint $i$ in the first frame of the input sequence, serving as the rest-pose reference for computing per-frame displacements. When orientation data is unavailable (e.g., for 2D pose estimations or position-only MoCap), we set $\mathbf{R}_{i}=\mathbf{I}_{3}$, reducing LBS to translation-only skinning—a key mechanism enabling format agnosticism.
Blending and SO(3) Projection. Each Gaussian $k$ is associated with the skeleton via blend weights $\mathbf{w}_{k}\in\Delta^{J-1}$ (on the probability simplex). These weights are pre-computed and fixed based on the skeleton’s topology: a joint Gaussian at joint $j$ receives a one-hot weight ($w_{k,j}{=}1$); a bone Gaussian interpolated between joints $a$ and $b$ at factor $\alpha$ has $w_{k,a}{=}1{-}\alpha,\;w_{k,b}{=}\alpha$ (zero elsewhere). Thus, the summation in Equation 3 reduces to at most two non-zero terms for any Gaussian. The blended translation and rotation are computed as:
@@
For each Gaussian $k$, we aggregate position $\mathbf{p}_{k}^{t}=\sum_{i}w_{k,i}\mathbf{j}_{i}^{t}$ and velocity $\mathbf{v}_{k}^{t}=\sum_{i}w_{k,i}\dot{\mathbf{j}}_{i}^{t}$ (finite-difference: $\dot{\mathbf{j}}_{i}^{t}=\mathbf{j}_{i}^{t}-\mathbf{j}_{i}^{t-1}$) from associated joints, concatenate with base appearance from a lightweight MLP, and process through a single-layer GRU for temporal modeling (it captures motion-phase cues such as acceleration vs. deceleration; see Appendix A.3 for ablations on temporal modeling choices). The NFM then predicts RGB and opacity residuals $(\Delta\text{RGB}_{k},\Delta\alpha_{k})$ plus a saliency gate $g_{k}$, which modulates the final opacity as $\alpha_{k}{=}\sigma(\alpha_{k}^{\text{base}}{+}\Delta\alpha_{k})\!\cdot\!\sigma(g_{k})$, allowing the network to suppress visually uninformative (e.g., stationary) primitives and amplify motion-salient ones. The final color blends the modulated appearance with a depth-based colormap: $\mathbf{C}_{k}=(1{-}\lambda)\sigma(\text{RGB}_{k}^{\text{base}}{+}\Delta\text{RGB}_{k})+\lambda\mathbf{C}_{k}^{\text{depth}}$, where $\sigma(\cdot)$ is the sigmoid function and $\lambda{=}\sigma(\theta_{\text{mix}})$ is a learnable mixing weight.
@@ 3.3 Vision-Language Backbone Integration
Generative classification via MQA. Since the MLLM is an auto-regressive generative model, we do not add a separate classification head. Instead, action recognition is cast as a multiple-choice question answering (MQA) task: the text prompt lists candidate action labels and instructs the model to generate a short answer ending with Label: <action>. The predicted label is extracted via pattern matching at inference. All training losses below are therefore standard auto-regressive cross-entropy on the target token sequence, with prompt tokens masked (see Appendix E.3.1 for prompt templates).
@@ 3.4 Cooperative Training Strategy
To address “brittle semantics” between similar actions (e.g., “rubbing hands” vs. “clapping”), we introduce a contrastive-like binary classification task. We construct hard negative pairs and ask the model: “Does this video show [Action A]?” (when it actually shows similar Action B). We update the renderer ($\Theta_{\text{render}}$) and projector ($\Theta_{\text{proj}}$); the model generates a single answer token (“Yes”/“No”) and we minimize the auto-regressive negative log-likelihood of this token—equivalently, binary cross-entropy on the positive-token probability—denoted $\mathcal{L}_{\text{Disc}}$ (see Appendix E.3.2). This directs the renderer to attend to subtle, discriminative motion details.
@@
Moving beyond recognition to deep understanding, we distill structured reasoning from a stronger teacher model (Hurst et al., 2024). We prompt the teacher to generate a causal chain of thought: step-by-step analysis of body-part movements, temporal evolution, and potential intent, followed by a conclusion. Our model is trained to generate this complete rationale given the rendered video. We optimize the renderer ($\Theta_{\text{render}}$), projector ($\Theta_{\text{proj}}$), and LLM (via LoRA, $\Theta_{\text{LoRA}}$) using auto-regressive loss $\mathcal{L}_{\text{CR}}$ over the full teacher token sequence (prompt tokens masked; see Appendix E.3.3). This instills structured, causal action understanding, enabling the model to explain why an action is classified as such.
@@
Finally, we freeze the mature renderer and update only the projector ($\Theta_{\text{proj}}$) and LLM ($\Theta_{\text{LoRA}}$) to refine the mapping for open-vocabulary recognition, minimizing auto-regressive cross-entropy loss $\mathcal{L}_{\text{MQA}}$ on the target label tokens (same as Stage 1).
@@
More design and implementation details on DrAction (Appendix E.2), Disc-FT and CR-Distill (Appendix E.3), as well as ablations on progressive training (Appendix A.1), are provided in the Appendix.
Table 1: Top-1 accuracy (%) of various methods on NTU-60 and NTU-120. Each split is denoted as X/Y, where X is the number of seen classes and Y is the number of unseen classes. The best results are in red, and the second-best are blue. Results for Qwen2.5-VL-7B and InternVL3-8B were obtained by rendering skeletons with the non-learnable 3D+Velocity renderer (same as in Table 4; $448{\times}448$) and finetuning on MQA for 6 epochs.
@@ 4.1 Experimental Setup
  • Finetuned MLLMs: Qwen2.5-VL and InternVL3 rendered with the non-learnable 3D+Velocity renderer (same as in Table 4; 12 frames, $448{\times}448$, identical front-view camera), finetuned on MQA recognition for 6 epochs.
@@ E.2.3 Neural Feature Modulator (NFM) Architecture
A single-layer GRU (hidden size 10) models temporal dependencies across frames, producing $\mathbf{h}_{k}^{t}=\text{GRU}(\mathbf{x}_{k}^{t},\mathbf{h}_{k}^{t-1})$. A two-layer MLP ($10\to 64\to 5$) with ReLU activation then predicts: